# Beyond Linearity and Time-homogeneity: Relational Hyper Event Models with Time-Varying Non-Linear Effects

### v2

@@ 4.2 Model fitting procedure
$f\left(\dot{\bm{x}}(t,I,J),\dot{t}\right)=\sum_{k=1}^{p=12}\sum_{q=1}^{Q}\left(\sum_{l=1}^{L}\alpha^{k}_{ql}a^{k}_{l}(\dot{t})\right)b^{k}_{q}\left(\dot{x}^{k}(t,I,J)\right)$
@@
Fitting the complete model, we observed extreme values for the effect of the covariate “Prior joint papers.” On closer inspection, only a very small number of non-events took a value different from zero (18 out of 1,416,353). This indicates that, for a randomly selected potential event, it is rare to find a pair of actors with previous co-authorship. In contrast, events often show nonzero values for this covariate. To avoid unstable non-linear estimates, we chose to include a linear effect for this covariate.
Excluded Covariate AIC Difference LogLik Difference Deviance Difference
Difference in prior papers -3886.01 1955.02 -3910.04
Author citation popularity -2858.46 1435.35 -2870.69
Difference in author citation popularity -408.09 215.52 -431.04
Paper outdegree popularity -433.07 232.89 -465.77
Prior papers -19288.83 9651.47 -19302.94
Prior joint papers (Linear) -3210100.35 1605046.84 -3210093.69
Collaborate with citing author -341.42 176.77 -353.54
Paper citation popularity -34990.31 17508.85 -35017.71
Paper pair cocitation -15622.95 7809.36 -15618.71
Author citation repetition -6249.01 3134.45 -6268.9
Cite paper and its references -19968.94 10001.37 -20002.73
Author self citation -33584.8 16789.89 -33579.78
Table 1: Contribution of individual network attributes. For each row in the table, we report the difference in AIC, log-likelihood (LogLik), and deviance between the full model (including all 12 covariates) and a reduced model that excludes the covariate listed in that row. This value reflects the difference in log-likelihood between the full model and the reduced model.
To perform model selection, we explore only a subset of the nested models. Namely, for each of the 12 presented covariates, we fitted a model including all the other covariates. We then assessed the effect of each covariate by comparing the change in AIC and log-likelihood between the full and partial models. These differences can be evaluated in Table 1. All twelve effects lead to an improvement in the AIC. Therefore, our final model contains all the mentioned covariates. To facilitate comparison with the baseline model assuming linear effects that are homogeneous over time, we fit standard RHEM Lerner et al. (2025) specified with the twelve effects; see Table 2.
@@
Difference in prior papers 0.85 0.03 30.84 0.00
Author citation popularity -1.14 0.03 -39.33 0.00
Difference in author citation popularity -0.70 0.03 -23.09 0.00
Paper outdegree popularity 0.32 0.01 24.49 0.00
Prior papers -0.42 0.03 -15.76 0.00
Prior joint papers 440.00 21.12 20.83 0.00
Collaborate with citing author 1.66 0.02 77.51 0.00
Paper citation popularity 4.55 0.03 162.28 0.00
Paper pair cocitation 40.58 1.37 29.56 0.00
Author citation repetition 38.74 1.63 23.80 0.00
Cite paper and its references 31.07 1.39 22.36 0.00
Author self citation 47.56 1.92 24.82 0.00
@@ 4.3 Evolving Dynamics of Scientific Collaboration and Impact
We first consider the interpretation of the covariate “Prior joint papers,” which was included with a linear effect. The estimated coefficient is 421.93, with a standard error of 26.99. This show a very strong, but not degenerate, effect of this variable. Furthermore, Table 1 highlights the importance of retaining this covariate in the model formulation.
@@
This covariate on a hyperedge $(I,J)$ at time $t$ is the average of the absolute pairwise differences in the number of previously published papers, taken over all pairs of authors in $I$. The baseline effect in Table 2 is positive, suggesting that teams of authors are typically composed of authors with large differences in the number of papers they have previously published, for example a PhD student together with her supervisor. The TVNLE plot in Figure 4 a) suggests a clearly non-monotone (hence non-linear) effect. If we fix a point on the horizontal axis (time) and go from bottom (covariate at its minimum value) to top (covariate at its maximal value), we find an increase in the event rate, up to about $x=1.5$, but then a decrease in the event rate. This means that the linear pattern (more difference in the number of prior papers implies a higher event rate) holds only for the lower values but gets reversed for higher values. The non-monotone pattern is qualitatively similar over time – although the maximum is less pronounced at the end of the observation period and sharper at the beginning.
@@
This covariate on a hyperedge $(I,J)$ at time $t$ is the average number of prior citations among all pairs of papers in $J$. The baseline effect in Table 2 is strongly positive, suggesting a tendency to partially copy the reference list of cited papers. However, the TVNLE plot in Fig. 4 b) shows that this very strong effect mostly holds around the end of the observation period, while earlier the effect is much weaker. Prior to 2011 the covariate even has a slightly non-monotone effect that has an increasing effect on the event rate for low values of the covariate and a decreasing effect for high values. The tendency to cite a paper together with some of its references is thus an example of a time-varying effect. The finding that the tendency to cite a paper together with some of its references becomes very strong only at the end of the observation period could be of substantive interest as it gives support to the interpretation that increasing use of publication databases and paper-search technology might strengthen this effect (Lerner et al., 2025). Since it becomes increasingly simple to follow paper citations forward and backward, and to retrieve the respective papers, the tendency to cite a paper and some of its references, or some of the papers previously citing it, may become stronger over time.
@@
This covariate on a hyperedge $(I,J)$ at time $t$ is the fraction of pairs $(i,j)\in I\times J$, such that $i$ is among the authors of $j$. The baseline effect in Table 2 is strongly positive, suggesting that authors have a tendency to cite their own prior work. The TVNLE plot in Fig. 5(a) suggests a more complex pattern where an approximately monotone increase can be found only at the beginning and end of the observation period, while for time in the interval $[2008,2016]$ the effect is clearly non-monotone, first increasing for low values of the covariate and then decreasing for high values. This covariate is thus an example of an empirical effect that is jointly time-varying and non-monotone.
@@
We have also found effects that are approximately monotone and nearly constant over time. The covariate “paper citation popularity” on a hyperedge $(I,J)$ at time $t$ is the average number of prior citations received by the papers in $J$. The baseline effect in Table 2 is positive, revealing the well-known “preferential attachment” or “rich-get-richer” effect in which papers that have been cited more often in the past are cited at a higher rate in the future. The TVNLE plot in Fig. 5(b) qualitatively confirms this by showing a monotone effect of the covariate throughout the observation period (although the increase is somewhat steeper in the middle and flatter for the earlier and later time points).
In summary, our empirical analysis suggests that in relevant hyperevent networks we can find effects that are approximately monotone and constant over time (paper citation popularity), time-varying but almost monotone (cite paper and its references), non-monotone but mostly constant over time (difference in prior papers), or jointly time-varying and non-monotone (self citation).
@@ 5 Discussion
In this paper, we have demonstrated how tensor product smooths can be used to model time-varying non-linear effects dynamics within the framework of relational hyperevent models. We applied this approach to study the dynamics of citation and impact. Particularly in the context of the time-sensitive relational event process, this is a valuable addition in the network modeling toolbox.
Strictly related to that, joint time-varying non-linear effects represent a powerful tool to discover patterns that not only may be non-linear but also non-monotonic. Indeed, when assuming linearity of the effect, we implicitly assume that either the effect is always increasing or always decreasing. This may not always be the case. As shown in the empirical application presented in this paper, there are situations where the effect of a covariate reaches a maximum or minimum and then reverses its trend. For example, certain factors may encourage the occurrence of events up to a certain threshold, but beyond that point, they may become counterproductive. A concrete case is author self-citation: citing a certain share of one’s own work seems natural and in fact may be nearly unavoidable, as the own prior work may be among the most related to a new publication. However, excessive self-citation could be perceived negatively by the research community and would not well connect the new work to that of others. These aspects would not be possible to capture in a purely linear setting.
We showed how the dynamics related to a specific driver of the hyperevent data-generating process can be represented using two-dimensional heatmaps. What is relevant about this approach is that each heatmap reflects the most appropriate functional form according to the data. As for time-varying and non-linear effects, when dealing with a joint time-varying non-linear effect, we let the contribution of each covariate take a data-driven shape. This increased flexibility comes at a computational cost. For each covariate, we need to estimate $L\times Q$, where $Q$ is the basis dimension related to non-linearity of the covariate ad $L$ is the basis dimension related to time. In the linear approach, for one covariate, we just need a single slope parameter. When dealing with many covariates, as in high-dimensional settings, this can pose a significant challenge — especially when estimation relies on inverting the Hessian matrix. It would therefore be important to adapt stochastic gradient descent techniques, such as those proposed in Filippi-Mazzola and Wit (2024b), to relational hyperevent models and, in particular, to these types of flexible effect specifications.
@@
Whereas in this paper we focused on relational hyperevent models, also in the usual REM context, when dealing with dyadic relational events, one could consider incorporating this more flexible type of effect. Modeling and inference guidelines would remain largely the same.
@@ 6 Acknowledgments
This work was supported by the Swiss National Science Foundation [grant 192549]; and the Deutsche Forschungsgemeinschaft (DFG) [Project No. 321869138].
@@ 8 Author contributions
MB: Conceptualization, Methodology, Software, Formal analysis, Writing - Original Draft; JL: Conceptualization, Methodology, Software, Resources, Data Curation, Writing - Review & Editing; EW: Conceptualization, Methodology, Writing - Review & Editing, Supervision;

### v3

@@ 5.2 Model fitting procedure
$f\left(\dot{\bm{x}}(t,I,J),\dot{t}\right)=\sum_{k=1}^{p=12}\sum_{q=1}^{Q}\left(\sum_{l=1}^{L}\alpha^{(k)}_{ql}a^{(k)}_{l}(\dot{t})\right)b^{(k)}_{q}\left(\dot{x}^{(k)}(t,I,J)\right)$
@@
Fitting the complete model, we observed extreme values for the effect of the covariate Prior joint papers. On closer inspection, only a very small number of non-events took a value different from zero (18 out of 1,416,353). This indicates that, for a randomly selected potential event, it is rare to find a pair of actors with previous co-authorship. In contrast, events often show nonzero values for this covariate. To avoid unstable non-linear estimates, we chose to include a linear effect for this covariate.
To perform model selection, we explore only a subset of the nested models. Namely, for each of the 12 presented covariates, we fitted a model including all the other covariates. We then assessed the effect of each covariate by comparing the change in Akaike Information Criterion (AIC) and log-likelihood between the full and partial models. These differences can be evaluated in Table 2. All twelve effects lead to an improvement in the AIC. Therefore, our final model contains all the mentioned covariates. To facilitate comparison with the baseline model assuming linear effects that are homogeneous over time, we fit standard RHEM lerner2024relational specified with the twelve effects; see Table 3.
@@
Difference in prior papers 0.85 0.03 30.84 <0.001
Author citation popularity -1.14 0.03 -39.33 <0.001
Difference in author citation popularity -0.70 0.03 -23.09 <0.001
Paper outdegree popularity 0.32 0.01 24.49 <0.001
Prior papers -0.42 0.03 -15.76 <0.001
Prior joint papers 440.00 21.12 20.83 <0.001
Collaborate with citing author 1.66 0.02 77.51 <0.001
Paper citation popularity 4.55 0.03 162.28 <0.001
Paper pair cocitation 40.58 1.37 29.56 <0.001
Author citation repetition 38.74 1.63 23.80 <0.001
Cite paper and its references 31.07 1.39 22.36 <0.001
Author self citation 47.56 1.92 24.82 <0.001
@@ 5.3 Evolving Dynamics of Scientific Collaboration and Impact
Prior joint papers.
We first consider the interpretation of the covariate prior joint papers, which, as mentioned above, was included with a linear effect to avoid unstable non-linear estimates. The estimated fixed time-inhomogeneous coefficient for this variable, when the others are modeled with TVNLE, is 421.93 (s.e. 26.99). It is important to emphasize that this coefficient refers to the transformed covariate, whose values are mapped to the $[0,1]$ interval. Nonetheless, the positive effect remains interpretable as such, given the monotonic nature of the transformation and the fact that the original covariate takes only positive values. Table 2 highlights the importance of retaining this covariate in the model formulation.
@@
This covariate on a hyperedge $(I,J)$ at time $t$ is the average of the absolute pairwise differences in the number of previously published papers, taken over all pairs of authors in $I$. The baseline effect in Table 3 is positive, suggesting that teams of authors are typically composed of authors with large differences in the number of papers they have previously published, for example a PhD student together with her supervisor. The TVNLE plot in Figure 9a suggests a clearly non-monotone (hence non-linear) effect. If we fix a point on the horizontal axis (time) and go from bottom (covariate at its minimum value) to top (covariate at its maximal value), we find an increase in the event rate, up to about $x=1.5$, but then a decrease in the event rate. This means that the linear pattern (more difference in the number of prior papers implies a higher event rate) holds only for the lower values but gets reversed for higher values. The non-monotone pattern is qualitatively similar over time – although the maximum is less pronounced at the end of the observation period and sharper at the beginning.
The positive baseline effect reported in Table 3, assuming a linear and homogeneous over time, suggests that scientists are more likely to coauthor a paper when potential team members differ more in their number of prior publications. In contrast, the non-linear effect in Figure 9 a) indicates that there is an “optimal” level of variation in prior publications that maximizes this covariate’s contribution to the log-rate of coauthoring a new paper; beyond this point, the coauthoring rate declines – here, “optimal” refers only to the co-authoring rate, not to paper quality.
@@
This covariate on a hyperedge $(I,J)$ at time $t$ is the average number of prior citations among all pairs of papers in $J$. The baseline effect in Table 3 is strongly positive, suggesting a tendency to partially copy the reference list of cited papers. However, the TVNLE plot in Fig. 9b shows that this very strong effect mostly holds around the end of the observation period, while earlier the effect is much weaker. Prior to 2011 the covariate even has a slightly non-monotone effect that has an increasing effect on the event rate for low values of the covariate and a decreasing effect for high values. The tendency to cite a paper together with some of its references is thus an example of a time-varying effect.
The finding that the tendency to cite a paper together with some of its references becomes very strong only at the end of the observation period could be of substantive interest as it gives support to the interpretation that increasing use of publication databases and paper-search technology might strengthen this effect (lerner2024relational). Since it becomes increasingly simple to follow paper citations forward and backward, and to retrieve the respective papers, the tendency to cite a paper and some of its references, or some of the papers previously citing it, may become stronger over time.
@@
This covariate on a hyperedge $(I,J)$ at time $t$ is the fraction of pairs $(i,j)\in I\times J$, such that $i$ is among the authors of $j$. The baseline effect in Table 3 is strongly positive, suggesting that authors have a tendency to cite their own prior work. The TVNLE plot in Fig. 10a suggests a more complex pattern where an approximately monotone increase can be found only at the beginning and end of the observation period, while for time in the interval $[2008,2016]$ the effect is clearly non-monotone, first increasing for low values of the covariate and then decreasing for high values. This covariate is thus an example of an empirical effect that is jointly time-varying and non-monotone.
The non-monotone effect in the time interval $[2008,2016]$ suggests that during this time period possible reference lists with approximately 5% to 7% of self-citations have the highest probability to be the reference lists of actual publication events, rather than sampled reference lists of non-events. In contrast publications with more than 7% self-citations become less likely during the time interval $[2008,2016]$. After 2016 the pattern becomes monotonically increasing, suggesting that in this period reference lists with higher ratios of self-citations are becoming increasingly more likely.
@@
We have also found effects that are approximately monotone and nearly constant over time. The covariate Paper citation popularity on a hyperedge $(I,J)$ at time $t$ is the average number of prior citations received by the papers in $J$. The baseline effect in Table 3 is positive, revealing the well-known “preferential attachment” or “rich-get-richer” effect in which papers that have been cited more often in the past are cited at a higher rate in the future. The TVNLE plot in Fig. 10b qualitatively confirms this by showing a monotone effect of the covariate throughout the observation period (although the increase is somewhat steeper in the middle and flatter for the earlier and later time points). The number of prior citations seems to monotonically increase a paper’s likelihood to be cited in the future and that this pattern does hardly change over time.
In summary, our empirical analysis suggests that in relational hyperevent networks we can find effects that are approximately monotone and constant over time (paper citation popularity), time-varying but almost monotone (cite paper and its references), non-monotone but mostly constant over time (difference in prior papers), and jointly time-varying and non-monotone (self-citation).
@@ 6 Discussion
Whereas this paper focused on relational hyperevent models, also in the usual REM context, when dealing with dyadic relational events, one could consider incorporating jointly time-varying non-linear effects. Modeling and inference guidelines would remain largely the same.
From a technical perspective, we have shown how tensor product smooths can be used to model time-varying non-linear effects dynamics within the framework of relational hyperevent models. We applied this approach to study the dynamics of citation and impact. Particularly in the context of the time-sensitive relational event process, this is a valuable addition in the network modeling toolbox.
Jointly time-varying and non-linear effects are a flexible tool for discovering patterns that not only may be non-linear but also non-monotonic. Indeed, when assuming linearity of the effect, we implicitly assume that either the effect is always increasing or always decreasing. As shown in the empirical application presented in this paper, there are situations where the effect of a covariate reaches a maximum or minimum and then reverses its trend. For example, certain factors may encourage the occurrence of events up to a certain threshold, but beyond that point, they may become counterproductive. A concrete case of non-monotonicity is author self-citation between 2008 and 2016: citing a certain share of one’s own work seems natural and in fact may be nearly unavoidable, as one’s own prior work may be among the most related to a new publication. However, excessive self-citation is perceived negatively by the research community and does not integrate the new work to that of others.
Certain patterns characterizing scientific collaboration and citations are evolving over time. In particular, several mechanisms exhibit changes in their behaviour towards the end of the observational window, such as self-citations and the tendency to cite chunks of the reference lists of cited works. Other mechanisms, such as collaboration with a citing author and author-citation repetition, also show changes in recent years, either in trend or intensity, compared to earlier periods. Another notable example, described in the supplementary materials, concerns the impact of an author’s prior publications. In the most recent period, the effect appears monotonic: the greater the number of prior publications, the more likely the author is to participate in a new publication. This contrasts with earlier periods, when teams with, on average, close-to-zero prior papers had a higher chance of publishing than those with a small but nonzero publication history. After reaching a minimum, the pattern then increases, reflecting the principle that it more likely to observe publications made by authors with a more extensive publication record.
Both these aspects, i.e. non-monotonicity as well as the temporal heterogeneity of an effect, would not be possible to capture with a purely linear model. Other approaches can be used to study the time-evolution of the impact of drivers on event rates. However, unlike the method proposed by kamalabad2023point, we do not need to explicitly test for change points, as such variations are captured by the smooth function of time included in each effect. However, our approach assumes that the time function is smooth, so sudden changes in the effects – permissible under the change point detection method – are not easily accommodated in our approach. This is generally not an issue for publication mechanisms, particularly given the yearly granularity of the data. uzaheta2024modeling considered the use of time-varying coefficients due to changes in network state in the context of DyNAMs. As they require the assumption of piecewise constant inter-arrival times, the rate of state switching (causing a change in parameters) must be lower than the rate of relational events. This can be a limiting assumption, particularly in the presence of time-varying exogenous effects, especially if unmeasured.
Although we propose tensor product smoother, other approaches exist for fitting jointly time-varying non-linear effects, e.g., using an isotropic smooth function of time and the covariate. However, isotropic smooths can perform poorly when the covariates are on different scales (wood2006low). Although we transformed both time and the covariate to the interval $[0,1]$ to mitigate such problems, the employed tensor product smooth approach is generally safer, especially when applying the model directly to data without intermediate transformations.
If prior knowledge is available about the functional form of an effect, the estimation method can be adapted using shape constraints. SCAMs (pya2015shape) extend GAMs by allowing component smooth functions to satisfy monotonicity or convexity constraints. This is achieved through a modified spline basis in which the constraints are enforced via a reparameterization of the spline coefficients.
We showed how the dynamics related to specific drivers of the hyperevent data-generating process can be represented using two-dimensional heatmaps. What is relevant about this approach is that each heatmap reflects the most appropriate functional form according to the data. As for time-varying and non-linear effects, when dealing with a jointly time-varying non-linear effect, we let the contribution of each covariate take a data-driven shape. This increased flexibility comes at a computational cost. For each covariate, we need to estimate $L\times Q$, where $Q$ is the basis dimension related to non-linearity of the covariate and $L$ is the basis dimension related to time. In the linear approach, for one covariate, we just need a single slope parameter. When dealing with many covariates, as in high-dimensional settings, this can pose a significant challenge – especially when estimation relies on inverting the Hessian matrix. It would therefore be important to adapt stochastic gradient descent techniques, such as those proposed in filippi2024stochastic, to relational hyperevent models and, in particular, to these types of flexible effect specifications.
Although we employ relational hyperevent modelling to study the dynamics of publication events, the recorded publication date does not necessarily reflect when the underlying collaboration actually occurred, nor when authors became aware of relevant prior work. Substantial delays may arise for many reasons, including journal publication frequency, editorial and reviewer availability, and the number and duration of review rounds. Moreover, in more recent times researchers often learn about new work through online pre-prints well before formal publication dates. As a result, recorded publication dates may be delayed relative to the true timing of scientific activity. Because relational hyperevent models explain publication dynamics as a function of previous collaborations and state of the field, such timing discrepancies may have practical implications. In particular, delayed publication can make past events appear less temporally proximate than they actually were, weakening recency effects in the observed data. Delays may also reorder the hyperevents themselves. The yearly time granularity of the data – although it may render several events as simultaneous when they are not – helps mitigating some of the issues due to the delay.
@@ 7 Acknowledgments
This work was supported by the Swiss National Science Foundation [grant 192549]; and the Deutsche Forschungsgemeinschaft (DFG) [Project No. 555455503].
@@ 9 Author contributions
MB: Conceptualization, Methodology, Software, Formal analysis, Writing - Original Draft; JL: Conceptualization, Methodology, Software, Resources, Data Curation, Writing - Review & Editing, Funding acquisition; EW: Conceptualization, Methodology, Writing - Review & Editing, Funding acquisition, Supervision;